OpenStudy (anonymous):

Simplify: x^-1/7*x^-1/6

OpenStudy (anonymous):

When you have the same base, you can just add the exponents. To add these ones, you need to put them over a common denominator.

OpenStudy (anonymous):

Okay did that :)

OpenStudy (jdoe0001):

\(\bf \Large x^{-\frac{1}{7}}\cdot x^{-\frac{1}{6}}\implies x^{-\frac{1}{7}+-\frac{1}{6}\implies \frac{?}{LCD?}}\)

OpenStudy (anonymous):

Okay, I think the answer is

OpenStudy (anonymous):

Is that correct? (:

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Awesome thanks! :) :)

OpenStudy (anonymous):

yup is right

OpenStudy (anonymous):

thanks for the help peoples!

OpenStudy (jdoe0001):

\(\bf \Large {x^{-\frac{1}{7}}\cdot x^{-\frac{1}{6}}\implies x^{-\frac{1}{7}+-\frac{1}{6}\implies \frac{-13}{42}}\implies \cfrac{1}{x^{\frac{13}{42}}}}\)

OpenStudy (jdoe0001):

yw